# Fully Dynamic Matching and Ordered Ruzsa-Szemerédi Graphs

### v2

@@ 4.2 The Overall Algorithm via Certificates and Caching Them
First, we prove that this algorithm obtains a $\varepsilon$-approximation matching.
@@
Given a contracted graph $G^{\prime}$ with $n^{\prime}$ vertices that the maximum matching size in $G^{\prime}$ is $O(n^{\prime})$ in $t$ updates, Section 4.2 maintains an $(1-\varepsilon$)-approximate maximum matching for $G^{\prime}$ in the dynamic model.
@@
Note that we proved in Observation 9 that if the maximal matching size is of order $n^{\prime}$, then we do not need to output a matching in $\varepsilon^{2}n^{\prime}$ updates.
By Section 4.2, set $S_{1}$ contains edge insertions, and set $S_{2}$ contains all the matchings we found earlier in the algorithm.
Using Proposition 3, we are given constant subsets of vertices that contain matchings of size at least $\varepsilon n$, we can run Section 4.1 on these subsets to output a $\varepsilon$-approximate matching in given subset $U$, if there are no large enough activated matchings in set $S_{2}\cup S_{1}$. Given that we aim to find edge-disjoint matchings, we need to find a maximal matching in $S_{2}\cup S_{1}$ and remove its endpoints from $U$.
After finding a matching $M^{\prime}$ in $S_{2}\cup S_{1}$ of size $\varepsilon\mu$, we output this matching; otherwise, we deactivate all the endpoints in $M^{\prime}$. Note that we are certificated by Proposition 3 to have a matching of size $\mu$ in $U$. Since for maximal matching $M^{\prime}$ in $S_{1}\cup S_{2}$ we have $|M^{\prime}|<\varepsilon\mu$, then the maximum matching size in $S_{1}\cup S_{2}$ is at most of size $2\varepsilon\mu$. This certifies the existence of a matching of size $(1-2\varepsilon)\mu$ in the rest of $U$. Now, we run Section 4.1 on graph $U$ with endpoints of $M^{\prime}$ removed with parameter $\varepsilon/(1-2\varepsilon)$. This implies that we get a matching of size $\varepsilon\mu$. Since we deactivate vertices of $S_{2}$ we only add edge-disjoint matchings to set $S_{2}$.
Note that any vertex in $G^{\prime}$ represents a bucket of vertices of $G$. By maintaining the number of vertices between two buckets in $G$, we can observe whether there is an edge between two vertices in $G^{\prime}$. This is how we can certify if a matching $M^{\prime}$ in $S_{2}$ is in $G^{\prime}$.
We proved that Section 4.2 outputs a $\varepsilon$-approximate matching $M_{U}$ in each subset of $U$ presented by Proposition 3. Note that given $M_{U}$ for each subset $U$, Proposition 3 finds a $(1-\varepsilon)$-approximation matching in $G^{\prime}$. This is true by the assumption that the maximum matching size of $G^{\prime}$ is $O(n^{\prime})$. ∎
Now that we introduced an algorithm for each contraction, we explore how to build and maintain each contraction. To partition updates to the contraction graphs that we maintain we need to run Proposition 4 to approximate the matching size, and then use Proposition 6 to figure out the contracted graph preserving the maximum matching, which we will do after each $\varepsilon^{2}n^{\prime}$ updates, given that the last contracted graph had $n^{\prime}$ vertices. We introduce Section 4.2 that given updates calculates the matching size and finds a suitable contraction for the graph at each state. For each update, Section 4.2 calls Section 4.2 on all the contractions, however, it only outputs the matching build by the matching preserving contraction.
We will run $O(\log^{2}n)$ parallel versions of Section 4.2 each for a maintained contracted graphs that we maintain. More formally we introduce Section 4.2:

### v3

@@ 4.2 Our Dynamic Algorithm via Lemma 6
4.3 Correctness of Section 4.2
In this section, we prove Lemma 9.
Proof of Lemma 9.
The algorithm works as follows. $c_{\text{updates}}$ is a counter for the number of updates. Let a phase be $t$ consecutive updates starting with $c_{\text{updates}}=0$. The algorithm maintains the fully dynamic graph $G$. All the insertions throughout a phase are added in $G_{\text{add}}$ (if an inserted edge is deleted, we delete it from $G_{\text{add}}$ too). All the deletions throughout a phase are added to $G_{\text{del}}$ (if a deleted edge is later inserted, we delete it from $G_{\text{del}}$ too).
First, we prove that MatchAndCertify outputs a $(1,\varepsilon n)$-approximate matching of graph $G$. To do this, we prove that the union of $E_{\text{dense}}=G-G_{\text{add}}$ and $E_{\text{sparse}}=G_{\text{add}}\cup H_{\text{cert}}-G_{\text{del}}$ is exactly $G$. To see this, note that
$E_{\text{dense}}\cup E_{\text{dense}}=(G-G_{\text{add}})\cup(G_{\text{add}}\cup H_{\text{cert}}-G_{\text{del}})=G\cup(H_{\text{cert}}-G_{\text{del}})=G.$
The second inequality holds because $G_{\text{add}}\cap G_{\text{del}}=\emptyset$. The last equality holds because any edge $e\in H_{\text{cert}}$ either belongs to $G$ or is deleted at some point, which means $H_{\text{cert}}-G_{\text{del}}\subseteq G$. By Lemma 6, this implies that MatchAndCertify outputs a $(1,\varepsilon n/2)$-approximate matching $M$ for $G$. Once $M$ is computed, we do not change $M$ within the next $\varepsilon n/2$ updates unless by applying the deletions on it. Each such deletion reduces the size of $M$ by at most 1. Therefore, after $\varepsilon n/2$ updates, we have $|M|\geq\mu(G)-\varepsilon n$, thus $M$ remains a $(1,\varepsilon n)$-approximate matching for $G$ throughout. ∎
4.4 Runtime Analysis of Theorem 1
Our goal in this section is to prove Lemma 10. Namely, we want to bound the running time of Section 4.2 by relating it to density of ORS graphs. The following Lemma 11 is our main tool for this connection to ORS graphs.
@@ 4.4 Runtime Analysis of Theorem 1
Let $G=(V,E=M_{1}\cup\ldots\cup M_{t})$ be an $n$-vertex graph where $M_{1},\ldots,M_{t}$ are edge-disjoint matchings of size at least $r$ each. For any $i\in[t]$, define $G_{i}=(V,E_{1}\cup\ldots\cup E_{i})$ and let $d_{i}$ be $|E[G_{i}]\cap V^{2}(M_{i})|/|M_{i}|$. Then for any $\alpha\in(0,1)$, it holds that
$\sum_{i=1}^{t}\frac{1}{d_{i}}=O\left(\frac{1}{\alpha^{2}}\cdot\mathsf{ORS}_{n}(r(1-\alpha))\cdot\log n\right).$
Let us first prove Lemma 10 via Lemma 11. The proof of Lemma 11 is presented afterwards.
Proof of Lemma 10.
All the steps of Section 4.2 can be implemented in $\widetilde{O}(1)$ time per update except for Section 4.2 where we spend $\widetilde{O}(t)$ time and Section 4.2 where we call MatchAndCertify. Since we call Section 4.2 every $t$ updates, the amortized cost of this line is still $\widetilde{O}(1)$. It remains to analyze the cost of Section 4.2. Note that Section 4.2 is called every $\varepsilon n/2$ updates, so its cost will have to be amortized over these $\varepsilon n/2$ updates.
Let us first discuss the time needed to prepare the inputs to MatchAndCertify. This input consists of graphs $E_{\text{dense}}=G-G_{\text{add}}$ and $E_{\text{sparse}}=G_{\text{add}}\cup H_{\text{cert}}-G_{\text{del}}$. For $E_{\text{dense}}$ we only need to provide adjacency matrix query access. Since each such query can be answered by making one adjacency matrix query to $G$ and one to $G_{\text{add}}$, we do not need to spend any time preparing $E_{\text{dense}}$. For $E_{\text{sparse}}$, we go over the edges of $G_{\text{add}},H_{\text{cert}},G_{\text{del}}$ one by one, but since these graphs are reset after $t$ updates, they only contain at most $t$ edges and the time spent is $O(t)$ which amortized over $\varepsilon n$ updates adds up to $O_{\varepsilon}(t/n)$.
Now, let us compute the running time of the algorithm based on the total calls to the MatchAndCertify. Each call to this algorithm, by Lemma 6 satisfies one of the two conditions $(C1)$ and $(C2)$. Each call that satisfies $(C1)$ runs in $\widetilde{O}_{\varepsilon}(n+|E_{\text{sparse}}|)$ time. Each call that satisfies $(C2)$ runs in $\widetilde{O}_{\varepsilon}(|E_{\text{sparse}}|+n^{2+\varepsilon}/d)$ time. Let us for now ignore the second term in the latter running time. The rest of the computation, amortized over $\varepsilon n$ updates, sums up to $\widetilde{O}_{\varepsilon}(|E_{\text{sparse}}|/n)=\widetilde{O}_{\varepsilon}(t/n)$.
Let us define a phase of the algorithm to be $t$ consecutive updates while $c_{\text{updates}}<t$. For analyzing the second term in the running time of the calls to MatchAndCertify that satisfy condition $(C2)$ of Lemma 6, we analyze their total cost over the whole phase and amortize this over $t$. Note that this is very different from our analysis before that amortizes the cost over a shorter sequence of $\varepsilon n$ updates.
Take the $i$-th call to Lemma 6 within the phase that satisfies $(C2)$ and let $M_{C_{i}}$ be its certificate, which is a matching in $E_{\text{dense}}$ and let $d_{i}:=|E_{\text{dense}}\cap V^{2}(M_{C_{i}})|/|M_{C_{i}}|$ be the average degree of $M_{C_{i}}$ in $E_{\text{dense}}$ the moment that it is returned. Note that the total cost the part of the algorithm that we ignored above can be written as
$\sum_{i}\widetilde{O}_{\varepsilon}(n^{2+\varepsilon}/d_{i})=\widetilde{O}_{\varepsilon}(n^{2+\varepsilon})\cdot\sum_{i}1/d_{i}.$ (1)
In what follows, we upper bound the sum $\sum_{i}1/d_{i}$.
Note that the set $E_{\text{dense}}$ used in the definition of $d_{i}$ does not remain the same throughout a phase, and particularly differs for each $d_{i}$. But here is the crucial observation. Since $E_{\text{dense}}=G-G_{\text{add}}$, the set $E_{\text{dense}}$ is decremental throughout the phase. This is because every edge inserted is inserted to both $G$ and $G_{\text{add}}$, thus $G-G_{\text{add}}$ remains unchanged, while deletions from $G$ will be deleted from $E_{\text{dense}}$ as well.
Now define $H_{\geq i}:=M_{C_{i}}\cup M_{C_{i+1}}\cup\ldots\cup M_{C_{k}}$. From our earlier discussion that the $M_{C_{i}}$’s are matchings in a decremental graph, all edges in $H_{\geq i}$ were present when certificate $M_{C_{i}}$ was outputted by Lemma 6. Hence, for the average degree $d^{\prime}_{\geq i}$ of $M_{C_{i}}$ in $H_{\geq i}$ we have
$d^{\prime}_{\geq i}:=\frac{|E(H_{\geq i})\cap V(M_{C_{i}})^{2}|}{|M_{C_{i}}|}\leq\frac{|E_{\text{dense}}\cap V(M_{C_{i}})^{2}|}{|M_{C_{i}}|}=d_{i},$ (2)
where here $E_{\text{dense}}$ is the input to Lemma 6 at the step where certificate $M_{C_{i}}$ was returned. Therefore, applying Lemma 11 on the sequence $M_{C_{k}},M_{C_{k-1}},\ldots,M_{C_{1}}$ (note the reverse order), we have
[…]
It should be noted that our Lemma 11 requires the matchings in the sequence to be edge-disjoint. This is indeed satisfied by our algorithm, since once a certificate $M_{C_{i}}$ is returned, we add its edges to $H_{\text{cert}}$ and these edges will never belong to $E_{\text{dense}}$ for the rest of the phase.
Plugged into 1, this bounds the total (remaining) running time of a phase by
$n^{2+\varepsilon}\cdot\widetilde{O}(\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n)).$
Amortized over $t$, the number of updates of the phase, this sums up to
$\frac{n^{2+\varepsilon}\cdot\widetilde{O}(\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}{t}.$
Combined with the earlier bound, we arrive at the claimed amortized update time of
$\widetilde{O}_{\varepsilon}\left(\frac{t}{n}+\frac{n^{2+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}{t}\right).\qed$
We now turn to prove Lemma 11. First we prove the following auxiliary claim that helps with the proof of Lemma 11.
Claim 12.
Let $G=(V,E=M_{1}\cup\ldots\cup M_{t})$ be an $n$-vertex graph where $M_{1},\ldots,M_{t}$ are edge-disjoint matchings of size at least $r$ each. For any $i\in[t]$, define $G_{i}=(V,E_{1}\cup\ldots\cup E_{i})$ and let $d_{i}$ be $|E[G_{i}]\cap V^{2}(M_{i})|/|M_{i}|$. Let $d$ be an upper bound on $d_{i}$ for all $i\in[t]$. Then for any $\alpha\in(0,1)$, it holds that
$t=O\left(\frac{d}{\alpha^{2}}\cdot\mathsf{ORS}_{n}(r(1-\alpha))\right).$
@@
Let $\varepsilon=\alpha/3$ and $\delta=\varepsilon^{2}$. Let $H=M_{1}\cup...\cup M_{t}$. Consider the following process for constructing a graph $H_{3}$, that we eventually show is an appropriate ORS graph.
  • Step 1. Let us define $x_{e}$ for any edge $e=(u,v)$ appearing in matching $M_{i}$ as follows
    $x_{e}:=\sum_{j>i:u,v\in V^{2}(M_{j})}\frac{1}{d_{j}}.$
    Define $B_{1}=\{e\mid x_{e}>1/\varepsilon\}$.
    Let $H_{1}\subseteq H$ include a matching $M_{i}\in H$ iff no more than $2\varepsilon r$ edges of $M_{i}$ belong to $B_{1}$. For each matching $M_{i}\in H_{1}$, we then set $M_{i}\leftarrow M_{i}-B_{1}$. Observe that $|M_{i}|\geq(1-2\varepsilon)r$ for any $M_{i}\in H_{1}$.
  • Step 2.
    Let $H_{2}\subseteq H_{1}$ include every matching $M_{i}\in H_{1}$ independently with probability $\frac{\delta}{2d_{i}}$.
  • Step 3. For any $i$, call an edge $e\in M_{i}\in H_{2}$ bad if
    $e\in\bigcup_{j>i:M_{j}\in H_{2}}V^{2}(M_{j}).$
    Let $B_{2}$ be the set of bad edges. Let $H_{3}\subseteq H_{2}$ include matching $M_{i}\in H_{2}$ iff there are at least $(1-\varepsilon)|M_{i}|$ edges in $M_{i}$ that do not belong to $B_{2}$. We then set $M_{i}\leftarrow M_{i}-B_{2}$ for any $M_{i}\in H_{3}$. Observe that $|M_{i}|\geq(1-\varepsilon)(1-2\varepsilon)r\geq(1-3\varepsilon)r$ for any $M_{i}\in H_{3}$.
From the construction above, it can be verified that $H_{3}$ is an ORS graph on matchings of size at least $(1-3\varepsilon)r$. What remains, is to lower bound the number of matchings that survive in $H_{3}$. This is what the rest of the proof focuses on.
First, we prove that $|B_{1}|\leq\varepsilon|E|$. Suppose for the sake of contradiction that $|B_{1}|>\varepsilon|E|$; then
$\sum_{e\in B_{1}}x_{e}>|B_{1}|/\varepsilon>|E|.$
On the other hand, we have
$\sum_{e\in B_{1}}x_{e}\leq\sum_{e\in E}x_{e}=\sum_{i\in[t]}\frac{|E[G_{i}]\cap V^{2}(M_{i})|}{d}=\sum_{i\in[t]}\frac{d_{i}|M_{i}|}{d}\leq\sum_{i\in[t]}|M_{i}|=|E|,$
which contradicts the inequality above. Here, the first equality follows by double counting: say matching $M_{i}$ “contributes” a value of $1/d$ to edge $e$ if $e\in E[G_{i}]\cap V^{2}(M_{i})$; then $x_{e}$ can be verified to be the total contribution of all matchings to $e$. The second equality follows from the definition of $d_{i}$. The inequality after that follows from $d_{i}\leq d$, which is by the assumption of the lemma.
Now that we know $|B_{1}|\leq\varepsilon|E|$, we get that
$|H_{1}|\geq t/2.$ (3)
Otherwise, more than $t/2$ matchings have more than $2\varepsilon r$ edges in $B_{1}$, meaning that $|B_{1}|>\frac{t}{2}\cdot 2\varepsilon r=\varepsilon rt=\varepsilon|E|$, a contradiction.
Let us now fix a matching $M_{i}\in H_{1}$ and fix an edge $e=(u,v)\in M_{i}$. We have
$\Pr[e\in B_{2}\mid M_{i}\in H_{2}]\leq\sum_{j>i:u,v\in V^{2}(M_{j})}\frac{\delta}{2d_{j}}=\delta x_{e}/2\leq\frac{\delta}{2\varepsilon}=\varepsilon/2,$
where the last inequality follows because $x_{e}\leq 1/\varepsilon$ for every edge in $M_{i}\in H_{1}$. By linearity of expectation, we have
$\operatorname{\textbf{E}}[|M_{i}\cap B_{2}|\mid M_{i}\in H_{2}]=\sum_{e\in M_{i}}\Pr[e\in B_{2}\mid M_{i}\in H_{2}]\leq\frac{\varepsilon}{2}|M_{i}|.$
Applying Markov’s inequality, we thus have
$\Pr\big{[}|M_{i}\cap B_{2}|\geq\varepsilon|M_{i}|\,\,\,\big{|}\,\,\,M_{i}\in H_{2}\big{]}\leq 1/2.$
Therefore,
$\Pr[M_{i}\not\in H_{3}\mid M_{i}\in H_{2}]\leq 1/2.$
This implies that